# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ Abstract.
We express parallel transport for several common matrix Lie groups with a family of pseudo-Riemannian metrics in terms of matrix exponential and exponential actions. The expression for parallel transport is preserved by taking the quotient under certain scenarios. In particular, for a Stiefel manifold of orthogonal matrices of size $n\times d$, we give an expression for parallel transport along a geodesic from time zero to $t$, that could be computed with time complexity of $O(nd^{2})$ for small $t$, and of $O(td^{3})$ for large $t$, contributing a step in a long-standing open problem in matrix manifolds. A similar result holds for flag manifolds with the canonical metric. We also show the parallel transport formulas for the generalized linear group, and the special orthogonal group under these metrics.
@@
Parallel transport, Stiefel manifolds, Flag manifolds, Matrix Exponential, Matrix manifolds, Reductive Group.
@@ 1. Introduction
In recent years, Riemannian geometry has found many applications in optimization, image, vision, and in modeling data with constraints and symmetry. A fundamental concept in geometry is parallel transport. While a tangent vector in Euclidean space can be transported between different points with almost no additional computation, moving a tangent vector between points in Riemannian geometry is a challenging yet fundamental problem. Parallel transport allows one to move tangent vectors along curves in a manifold in a way consistent with geometry, which is important in clustering problems, and in algorithms that require accounting for past increments. For example, the well-known optimization methods L-BFGS or conjugate gradient on an Euclidean space keep a history of past movements and use them to decide the next moves. These increments are represented as tangent vectors at different points in the iteration and must be transported to the current iterative point for aggregation. Similar issues arise in computer vision.
In the now classical paper [5] twenty-five years ago, the authors twice mentioned that there was no known $O(nd^{2})$ algorithm for parallel transport on the Stiefel manifold $\mathrm{St}_{n,d}$. To our best knowledge, this problem remains unsolved. In this paper, we contribute toward a solution for this problem with the help of exponential action. We show with an appropriate transformation, a class of metrics on matrix manifolds have geodesics computable by matrix exponential, while the parallel transport equation could be reduced to a linear ODE with constant coefficients, and thus could be solved by exponential action. For the case of the Stiefel manifolds, we show any tangent vector could be split into two parts, one in a submanifold similar to the great circle, where we can use exponential action. The other part, normal to a vector subspace containing the geodesic, can be transported using matrix exponential. We establish a similar result for flag manifolds with the canonical metric. We also show that the formula for the Stiefel manifold is a special case for pseudo-Riemannian metrics in a family with closed-form geodesic and parallel transport. Besides the state-of-the-art algorithm in the Stiefel case, we hope our results will provide insight into the properties of parallel transport.
@@ 1.1. Contribution
We give an explicit, with cost $O(C_{1}nd^{2}+C_{2}td^{3})$ for the parallel transport of Stiefel manifolds. Our formula for the parallel transport of a tangent vector $\eta$ stating at a point $Y\in\mathrm{St}_{n,d}$, the Stiefel manifold of orthogonal matrices in $\mathbb{R}^{n\times d}$ equipped with the metric eq. 44 defined by a parameter $\alpha>0$ (the canonical metric corresponds to $\alpha=\frac{1}{2}$ and the Euclidean embedded metric corresponds to $\alpha=1$) along a geodesic with initial speed $\xi$ is given by
(1) $\begin{gathered}\Delta(t)=\begin{bmatrix}Y|Q\end{bmatrix}\operatorname{expm}(t\begin{bmatrix}2\alpha A&-R^{\mathsf{T}}\\ R&0\end{bmatrix})\operatorname{expv}(t\mathrm{P},\begin{bmatrix}Y|Q\end{bmatrix}^{\mathsf{T}}\eta)\operatorname{expm}(t(1-2\alpha)A)\\ +(\eta-YY^{\mathsf{T}}\eta-QQ^{\mathsf{T}}\eta)\operatorname{expm}(t(1-\alpha)A)\end{gathered}$
where $Q$ is an orthogonal basis of the column vectors of $\eta-YY^{\mathsf{T}}\eta$, $A$ and $R$ are such that $\xi=YA+QR$ and $\mathrm{P}$ is an operator defined in eq. 52, the column-wise concatenation of $Y$ and $Q$ is denoted $\begin{bmatrix}Y|Q\end{bmatrix}$, and $\operatorname{expv}$ is the exponential action. The formula can be evaluated with cost $O(C_{1}nd^{2}+C_{2}td^{3})$ with $n>d$ because $Q,A,R$ could be computed with cost $O(nd^{2})$ [5, Corollary 2.2], the matrix exponentials is of cost $O(d^{3})$ (or more precisely $O(d^{3}\log_{2}t))$, the exponential action is of cost $O(td^{3})$ and the matrix multiplications are $O(nd^{2})$. Thus, the time complexity is $O(C_{1}nd^{2}+C_{2}td^{3})$. For small $t$, which is the situation in optimization problems, the first $O(nd^{2})$ term dominates the cost, whereas, for large $t$, the second $O(td^{3})$ will dominate.
@@
The framework to derive this formula is also applicable to manifolds related to Stiefel manifolds, giving us a new parallel transport formula for flag manifolds with the canonical metric and recovering the known formula for Grassmann manifolds.
We also show the expression of parallel transport in terms of exponential action is applicable to a larger class of matrix groups and certain quotient spaces. We denote the relevant subspaces transposable subspaces, a property that in the Lie group literature is denoted reductivity but that we here use in the matrix context to obtain closed-form geodesics and parallel transport. With this result, we show explicit parallel transport formulas for this family of metrics for the generalized linear groups, and the special orthogonal groups. While not discussed here, the symplectic Lie algebra [6] is transposable, thus the results here also apply to symplectic groups.
@@ 1.2. Related literature and State-of-the-Art
Closed-form geodesics formulas for the generalized linear group are in [28, 18, 3]. In the original paper [5], Edelman et. al. gave an equation (equation 2.27) for parallel transport for the Stiefel manifold with the canonical metric. The differential equation is linear but even with exponential action, the cost is $O(n^{3})$. They also mentioned that this formula is a special case for reductive spaces (in particular, compact groups) and also gave an $O(nd^{2})$ algorithm for the Grassmann transport. As mentioned, reductivity overlaps, if not identical with our notion of transposability, but the latter notion is easy to formulate and leads quickly to explicit and computationally efficient formulas. We avoid the $O(n^{3})$ cost by exploiting a special property of Stiefel and flag manifolds, including Grassmann manifolds.
In [12], the authors give a formula for parallel transport for Stiefel manifolds in the Euclidean embedded metric ($\alpha=1$ in eq. 1). The formula is in terms of a function $F$ satisfying a matrix ODE, depending on the equation of the curve, and requires vectorization. The geodesic is calculated with $O(n^{3})$ cost for a given time. Other than that, we are not aware of progress in parallel transport for Stiefel manifolds. Numerically, in [9, 16, 15], the authors introduce numerical algorithms to solve transport problems, in particular the ladder schemes. As closed-form formulas are not available most of the time, they are of great importance.
The family of metrics considered here could be considered a deformation of a pseudo-Riemannian metric (also called semi-Riemannian metric) to a Riemannian metric, see [8], the deformation is often called the Cheeger deformation, although it has also been used by other authors. The Cheeger deformation comes from a (pseudo-Riemannian) submersion, thus, in geometry, is used to produce metrics with desired curvature properties. Geodesics could be lifted in a submersion.
In [13], the authors analyze sub-Riemannian geodesic arising from this family and note that this type of deformation produces closed-form geodesic. The resulting induced metrics for Stiefel manifolds are studied in [11]. In [20], one of the authors compute the connection and curvature for this family in the case of a bi-invariant metric, and proved proposition 1 in that case. Flag manifolds are studied by many authors, where Riemannian Hessian, exponential map, logarithm and the Levi-Civita connection are studied in [30, 24, 27, 21, 19].
@@ 1.3. Outline
In the next section, we review the background required in the paper, including bilinear form and pseudo-Riemannian geometry. We introduce the concept of transposable Lie-subalgebra, which includes $\mathfrak{gl}$ and $\mathfrak{so}$. We introduce the family of tractable metrics in theorem 1 and discuss the condition for the metrics to be Riemmanian.
We show the geodesic and parallel transport formula for generalized linear groups in section 4 and special orthogonal groups in section 5. In section 6, we show the quotient by a subgroup of particular structures will give closed-form geodesics. The parallel transport equation can be reduced to a linear ODE with variable coefficients in the general case, and constant coefficients, solvable by exponential action in special cases. In section 7, we provide parallel transport formulas for the Stiefel and flag manifolds, with a discussion of Grassmann manifolds in appendix A. In section 9, we provide numerical experiments. We end the paper with a discussion.
@@ 2. Background Review
2.1. Exponential Action
The solution of a linear ODE with constant coefficient $\dot{X}=AX$ and initial condition $X(0)=X_{0}$ is given by $\operatorname{expm}(tA)X_{0}$, where typically $X$ is a vector and $A$ is a matrix, where $\operatorname{expm}$ is the matrix exponential. We often need to treat the case where $X_{0}$ is a matrix and $A$ is an operator on a matrix space. In that situation, which is the situation encountered in this article, computing the exponential by vectorization is usually not efficient. The better alternative is to compute the exponential action, denoted by $\operatorname{expv}(tA,X_{0})$. Instead of evaluating the full matrix exponential, we approximate the action $\operatorname{expm}(tA)X_{0}$ more efficiently, often [2] by evaluating a scaling and squaring of a Taylor expansion of the operator exponential operating on $X_{0}$.
Specifically, the popular library algorithm for exponential action in Python (function $expm{\_}multiply$ in the scipy package [29] and similar functions for other languages, $expmv$ in MATLAB and $expv$ in Julia) are based on the work of Al-Mohy and Higham [2], approximating $e^{A}B\approx(T_{m}(s^{-1}A))^{s}B$, where $T_{m}$ denotes the $m$ terms Taylor series expansion of $e^{A}$. The cited paper provides an algorithm to find a pair of optimal $(m,s)$, based on estimates of $\|A^{p}\|^{1/p}_{1}$ for several values of $p$. The paper found the higher $p$ estimates are beneficial when $A$ is not normal, for example, when $A$ is upper triangular. These estimates are used to determine the Taylor order $m$.
In these implementations, the $1$-norm is estimated using the algorithm in [10]. We will use $expm{\_}multiply$ from scipy for $\operatorname{GL}(n)$ and $\mathrm{SO}(n)$, while we will show a customized implementation based on efficient estimates of the $1$-norm of the operator $\mathrm{P}$ for Stiefel and flag manifolds in section 9 could speed up the algorithm significantly, without the higher $p$ estimates.
@@ 2.2. Symmetric Bilinear forms
As we will discuss pseudo-Riemannian metric, it is appropriate to review some quick facts about bilinear forms, which could be considered as inner products without the positive-definite requirement. The main results that we need are about subspaces. A symmetric bilinear form, or just form on a vector space $\mathcal{E}$ is a scalar-valued bilinear map from $\mathcal{E}\times\mathcal{E}$ to $\mathbb{R}$, denoted by $\langle v,w\rangle$. For brevity, we will omit the word symmetric as we will only work with symmetric bilinear forms. If we identify $\mathcal{E}$ with $\mathbb{R}^{n}$ and represent the form by a matrix $\mathrm{B}$, then $\langle v,w\rangle=\langle v,w\rangle_{\mathrm{B}}=v^{\mathsf{T}}\mathrm{B}w$ for $v,w\in\mathcal{E}$. It is nondegenerate (or we also say $\mathcal{E}$ is nondegenerate) if $\mathrm{B}$ is nonsingular, or $\langle v,w\rangle=0$ for all $w\in\mathcal{E}$ implies $v=0$. This latter formulation is useful if $\mathcal{E}$ is a matrix space. We can restrict the form $\langle$,$\rangle$ , ⟩ to a subspace $\mathcal{V}\subset\mathcal{E}$. In contrast to the positive-definite case, a form may be degenerate on $\mathcal{V}$ even if it is nondegenate on $\mathcal{E}$. We define the orthogonal complement $\mathcal{V}_{{}_{\perp}}$ of $\mathcal{V}$ as follows
@@
Then $\mathcal{V}_{{}_{\perp}}$ may intersect $\mathcal{V}$ nontrivially. However, we have a generalization of several well-known results for inner product spaces below
@@
A subspace $\mathcal{V}\subset\mathcal{E}$ is nondegenerate if and only if $\mathcal{V}_{{}_{\perp}}\cap V=\{0\}$. In that case, we have a unique decomposition $\mathcal{E}=\mathcal{V}\oplus\mathcal{V}_{{}_{\perp}}$, or any vector $w\in\mathcal{E}$ could be written in a unique way as $w=w_{\mathcal{V}}+w_{\mathcal{V}_{\perp}}$, with $w_{\mathcal{V}}\in\mathcal{V},w_{\mathcal{V}_{\perp}}\in\mathcal{V}_{\perp}$. The map $\Pi_{\mathcal{V}}:w\mapsto w_{\mathcal{V}}$ is a projection, the orthogonal projection to $\mathcal{V}$. It satisfies
(3) $\langle\Pi_{\mathcal{V}}w,v\rangle=\langle w,v\rangle.$
Proof.
The equivalence of nondegeneracy of $\mathcal{V}$ and $\mathcal{V}_{{}_{\perp}}\cap V=\{0\}$ is just a rewording of the definitions. As $\mathcal{V}$ is nondegenerate, given a basis $\{v_{i}\}_{i=1}^{d}$, of $\mathcal{V}$, with $d=\dim\mathcal{V}$, let $C=C_{ij}=(\langle v_{i},v_{j}\rangle)_{i,j=1}^{d}$, and $v^{j}=\sum_{i}(C^{-1})_{ji}v_{i}$, then the set $\{v^{j}\}_{j=1}^{d}\subset\mathcal{V}$ satisfies $\langle v_{k},v^{j}\rangle=(C^{-1}C)_{jk}=\delta_{jk}$.
Define $w_{\mathcal{V}}=\sum_{i}\langle v^{i},w\rangle v_{i}\in\mathcal{V}$. Then we can check $\langle w_{\mathcal{V}},v^{j}\rangle=\sum_{i}\langle v^{i},w\rangle\delta_{ij}=\langle v^{j},w\rangle$. By linearity, $\langle w_{\mathcal{V}},v\rangle=\langle w,v\rangle$ for all $v\in\mathcal{V}$, or $\langle w-w_{\mathcal{V}},v\rangle=0$, hence $w-w_{\mathcal{V}}\in\mathcal{V}_{{}_{\perp}}$, and $w=w_{\mathcal{V}}+(w-w_{\mathcal{V}})\in\mathcal{V}\oplus\mathcal{V}_{{}_{\perp}}$. ∎

### v2

@@ Abstract.
We express parallel transport for several common matrix Lie groups with a family of pseudo-Riemannian metrics in terms of matrix exponential and exponential actions. The metrics are constructed from a deformation of a bi-invariant metric and are naturally reductive. There is a similar picture for homogeneous spaces when taking quotients satisfying a general condition. In particular, for a Stiefel manifold of orthogonal matrices of size $n\times d$, we give an expression for parallel transport along a geodesic from time zero to $t$, that could be computed with time complexity of $O(nd^{2})$ for small $t$, and of $O(td^{3})$ for large $t$, contributing a step in a long-standing open problem in matrix manifolds. A similar result holds for flag manifolds with the canonical metric. We also show the parallel transport formulas for the general linear group and the special orthogonal group under these metrics.
@@
Parallel transport, Stiefel manifolds, Flag manifolds, Matrix Exponential, Exponential Action, Matrix manifolds, Reductive Group.
@@ 1. Introduction
In recent years, Riemannian geometry has found many applications in optimization, image, vision, and data modeling with constraints and symmetry. A fundamental concept in geometry is parallel transport. While a tangent vector in Euclidean space can be transported between different points with almost no additional calculation, moving a tangent vector between points on a Riemannian manifold is more challenging computationally. Parallel transport allows one to move tangent vectors along curves in a manifold in a way consistent with geometry, which is important in clustering problems, and in algorithms that require accounting for past increments. For example, the well-known optimization methods L-BFGS or conjugate gradient on an Euclidean space keep a history of past movements and use them to decide the next moves. These increments are represented as tangent vectors at different points in the iteration and must be transported to the current iterative point for aggregation. Similar issues arise in computer vision.
In the now classical paper [7] twenty-five years ago, the authors twice mentioned that there was no known $O(nd^{2})$ algorithm for parallel transport on the Stiefel manifold $\mathrm{St}_{n,d}$. To the best of our knowledge, this problem remains unsolved. In this paper, we contribute toward a solution for this problem with the help of exponential action. We show with an appropriate transformation, a class of metrics on matrix manifolds have geodesics computable by matrix exponential, while the parallel transport equation could be reduced to a linear ODE with constant coefficients, and thus could be solved by exponential action. For the case of the Stiefel manifolds, we show any tangent vector could be split into two parts, one in a submanifold similar to a great circle, where we can use exponential action. The remaining part, normal to a vector subspace containing the geodesic, can be transported using matrix exponential. We also show that the formula for the Stiefel manifold is a special case for pseudo-Riemannian metrics in a family with closed-form geodesic and parallel transport. Besides the state-of-the-art algorithm in the Stiefel case, we hope our results will provide insight into the properties of parallel transport.
@@ 1.1. Contribution
We give an explicit formula eq. 61 with evaluation cost $O(C_{1}nd^{2}+C_{2}td^{3})$ for the parallel transport along a geodesic segment of length $t$ on the Stiefel manifold $\mathrm{St}_{n,d}$ of orthogonal matrices of size $n\times d$ ($d<n$) eq. 61 for constants $C_{1}$ and $C_{2}$. The cost $O(C_{2}td^{3})$ comes from the evaluation of an exponential action. This time-dependent cost is less significant for a shorter time, which is the typical situation in optimization. Thus, in this case, the remaining cost $O(C_{1}nd^{2})$ becomes dominant. For a bounded $t$, the algorithm could also be considered of time complexity $O(nd^{2})$, and since Stiefel manifolds are compact, there is an upper bound to the distance between two points, thus, the algorithm is practically $O(nd^{2})$ for reasonable time tt [^1] We thank a reviewer for this observation.. While the exponential action could be expensive, we propose a custom implementation based on a norm estimate that speeds up the evaluation significantly. We also hope the closed-form formula provides new insights into parallel transports, leading to new transport algorithms and applications.
@@
The framework to derive this formula also applies to manifolds related to Stiefel manifolds, giving us a new parallel transport formula for flag manifolds with the canonical metric and recovering the known formula for Grassmann manifolds.
As explained, this result is based on the observation that transport in directions tangent to a Stiefel submanifold (similar to a great circle) could be done efficiently by solving an equivalent problem that is reduced in dimension and is called baby problem in [35, chapter 5], while transport in directions orthogonal to this submanifold could be done by matrix exponential. This observation may apply to other families of manifolds.
The baby problem is solved by lifting to a matrix Lie group. This solution generalizes to matrix Lie groups and homogeneous spaces equipped with a deformation metric [5, 18, 6]. In particular, given a Lie group $\mathcal{G}$ with a bi-invariant pseudo-Riemannian metric represented by a symmetric bilinear form on its Lie algebra $\mathfrak{g}$ and a Lie subalgebra $\mathfrak{a}\subset\mathfrak{g}$ such that the bilinear form is nondegenerate when restricted to $\mathfrak{a}$, if we replace the metric on $\mathfrak{a}$ with another bi-invariant metric then the resulting metric is naturally reductive. The geodesic equation and geodesics are explicit, while parallel transport could be computed by exponential action. The replacement metric on $\mathfrak{a}$ is often a scalar multiple of the original metric. In some examples, we can change the sign of the metric on $\mathfrak{a}$, modifying the bi-invariant pseudo-Riemannian metric on $\mathfrak{g}$ to a naturally reductive Riemannian one, as in the case of $\operatorname{GL}(n)$ in [42] and section 5.
For a homogeneous manifold $\mathcal{G}/\mathcal{K}$ with $\mathcal{G}$ equipped with a deformation metric, we give a general condition in theorem 3 for the manifold $\mathcal{G}/\mathcal{K}$ to be also naturally reductive, with geodesics and parallel transport computable by exponential action.
With this result, we show explicit parallel transport formulas for this family of metrics for the general linear groups and the special orthogonal groups. While not discussed here, the results also apply to symplectic groups [8]. In particular, the Euclidean embedded metric of Stiefel manifolds and $\operatorname{GL}(n)$ with the Frobenius metric [7, 42] are naturally reductive. While instances of this result are already known in the literature [6, 19, 16], the explicit ODE for parallel transport, based on [39, 16] and the observation that certain tractable Riemannian metrics on non compact groups could be considered a deformation of the pseudo-Riemannian trace form is novel.
The nondegeneracy of the trace form as a bi-invariant metric is guaranteed if we restrict to Lie subalgebras that are equal to their own transpose, such as $\mathfrak{so}(n)$ or the full $n\times n$ matrix algebra, as opposed to the algebra of upper triangular matrices.
@@ 1.2. Related literature and State-of-the-Art
Closed-form geodesics formulas for the general linear group are in [42, 24, 3]. In the original paper [7], Edelman et. al. gave an equation (equation 2.27) for parallel transport for the Stiefel manifold with the canonical metric. The differential equation is linear but even with exponential action, the cost is $O(n^{3})$. They also mentioned that this formula is a special case for reductive spaces and also gave an $O(nd^{2})$ algorithm for the Grassmann transport. The paper referenced the thesis [39], which gave a parallel transport formula for a quotient $\mathcal{G}/\mathcal{H}$ in its Proposition 2.12. The proof assumed a bi-invariant metric on $\mathcal{G}$. When the metric is constructed from a deformation, (see theorems 1 and 3), which is essentially a quotient metric on a larger group $\tilde{\mathcal{G}}$ with a bi-invariant metric, we can apply this result to compute the parallel transport. Exploiting a special property of Stiefel and flag manifolds to split the transported vector as noted previously, we avoid the $O(n^{3})$ cost.
In [17], the authors give a formula for parallel transport for Stiefel manifolds in the Euclidean embedded metric ($\alpha=1$ in eq. 54). The formula is in terms of a function $F$ satisfying a matrix ODE, depending on the equation of the curve, and requires vectorization. The geodesic is calculated with $O(n^{3})$ cost for a given time. Vector transport [1, chapter 8], often used in Riemannian optimization, is another method to transport tangent vectors on manifolds. They are often less expensive computationally, although they are generally not isometries (but there are remedies [38] in Riemannian optimization). We are not aware of other progress in parallel transport for Stiefel manifolds. Numerically, in [13, 22, 21], the authors introduce numerical algorithms to solve transport problems, in particular, the ladder schemes. As closed-form formulas are not available most of the time, they are of great importance.
The family of metrics considered here could be considered a deformation of a pseudo-Riemannian metric (also called semi-Riemannian metric) to a Riemannian metric, see [12], the deformation is often called the Cheeger deformation, although it has also been used by other authors [18, 37]. The deformation comes from a (pseudo-Riemannian) submersion and is often used to produce metrics with desired curvature properties. In [6, Theorem 1], the authors analyzed naturally reductive space arising from this construction for compact groups. Our theorem 1 could be considered a partial extension of their construction to the pseudo-Riemannian case. The article [11] considers the non-compact spaces from a Lie theoretic approach, which we complement by offering an explicit metric construction and geometric formulas.
In [19], the authors analyze sub-Riemannian geodesics arising from this family and note that this type of deformation produces closed-form geodesic. The resulting induced metrics for Stiefel manifolds are studied in [16]. In [30], one of the authors computed the connection and curvature for this family in the case of a Riemannian bi-invariant metric. Flag manifolds are studied by many authors, where Riemannian Hessian, exponential map, logarithm, and the Levi-Civita connection are studied in [44, 34, 40, 31, 29].
@@ 1.3. Outline
In the next section, we review the background required in the paper, including bilinear form and pseudo-Riemannian geometry. We introduce the concept of transposable Lie-subalgebra, which includes $\mathfrak{gl}$ and $\mathfrak{so}$. We introduce intertwining operators and the family of deformation metrics in theorem 1 and discuss the condition for the metrics to be Riemmanian.
We show the geodesic and parallel transport formula for general linear groups in section 5 and special orthogonal groups in section 6. In section 7, we show the quotient by a subgroup with a particular structure has closed-form geodesics. The parallel transport equation can be reduced to a linear ODE with variable coefficients in the general case, and constant coefficients, solvable by exponential action in special cases. In section 8 and section 9, we provide parallel transport formulas for the Stiefel and flag manifolds, with a discussion of Grassmann manifolds in appendix B. In section 10, we provide numerical experiments. We end the paper with a discussion. In appendix A we review the exponential action and prove our norm estimate.
1.4. Notations
The general linear group $\operatorname{GL}(n)$ is the group of all invertible matrices in $\mathbb{R}^{n\times n}$, the connected subgroup with positive determinant is $\operatorname{GL}^{+}(n)$, the corresponding Lie algebra $\mathfrak{gl}(n)$ could be identified with $\mathbb{R}^{n\times n}$. The special orthogonal group $\mathrm{SO}(n)$ is the group of orthogonal matrices of determinant $1$, its Lie algebra $\mathfrak{so}(n)$ could be identified as a vector space with the space of antisymmetric matrices $\mathrm{Skew}_{n}$. Its orthogonal complement is the space $\mathrm{Sym}_{n}$ of symmetric matrices. The Stiefel, flag, and Grassmann manifolds with notations $\mathrm{St}_{n,p},\mathrm{Gr}_{n,p},\operatorname{Flag}(\vec{d})$ will be introduced later. We write $\mathrm{sym}$ for the symmetrize operator $A_{\mathrm{sym}}=\frac{1}{2}(A+A^{\mathsf{T}})$ for $A\in\mathbb{R}^{n\times}$ and $\mathrm{skew}$ for the antisymmetrize operator $A_{\mathrm{skew}}=\frac{1}{2}(A-A^{\mathsf{T}})$. The notation $\mathcal{P}_{{}_{\mathcal{V}}}$ denotes the projection to a subspace $\mathcal{V}$.
In a matrix space $\mathbb{R}^{n\times n}$, we define the $\operatorname{ad}$ operator $\operatorname{ad}_{\mathtt{X}}\mathtt{Y}=[\mathtt{X},\mathtt{Y}]=\mathtt{X}\mathtt{Y}-\mathtt{Y}\mathtt{X}$ for two matrices $\mathtt{X},\mathtt{Y}$ and the $\operatorname{Ad}$ operator $\operatorname{Ad}_{K}\mathtt{X}=K\mathtt{X}K^{-1}$ for an invertible matrix $K\in\mathbb{R}^{n\times n}$.
@@ 2. Background Review
We treat exponential action in the appendix. The main indefinite linear algebra and differential geometry background are reviewed here.
@@ 2.1. Symmetric Bilinear forms
As we will discuss pseudo-Riemannian metric, it is appropriate to review some quick facts about bilinear forms, which are indefinite scalar products. While we work with the real case, the proofs and results in the main reference [10] for the Hermitian case carry through readily. The main results that we need are about subspaces. A symmetric bilinear form, or just form, or scalar product on a vector space $\mathcal{E}$ is a scalar-valued bilinear map from $\mathcal{E}\times\mathcal{E}$ to $\mathbb{R}$, denoted by $\langle v,w\rangle$. For brevity, we will omit the word symmetric as we will only work with symmetric bilinear forms. If we identify $\mathcal{E}$ with $\mathbb{R}^{n}$ and represent the form by a matrix $\mathcal{B}$, then $\langle v,w\rangle=\langle v,w\rangle_{\mathcal{B}}=v^{\mathsf{T}}\mathcal{B}w$ for $v,w\in\mathcal{E}$. It is nondegenerate (or we also say $\mathcal{E}$ is nondegenerate) if $\mathcal{B}$ is nonsingular, or $\langle v,w\rangle=0$ for all $w\in\mathcal{E}$ implies $v=0$. This latter formulation is useful if $\mathcal{E}$ is a matrix space. We can restrict the form $\langle,\rangle$ to a subspace $\mathcal{V}\subset\mathcal{E}$. In contrast to the positive-definite case, an indefinite scalar product may be degenerate on $\mathcal{V}$ even if it is nondegenerate on $\mathcal{E}$. We define the orthogonal complement $\mathcal{V}_{{}_{\perp}}$ of $\mathcal{V}$ as follows
@@
Then $\mathcal{V}_{{}_{\perp}}$ may intersect $\mathcal{V}$ non-trivially. The real case of [10, Proposition 2.2.2] generalizes several well-known results for inner product spaces in the lemma below.
@@
A subspace $\mathcal{V}\subset\mathcal{E}$ is nondegenerate if and only if $\mathcal{V}_{{}_{\perp}}\cap\mathcal{V}=\{0\}$. Assuming this condition, we have a unique decomposition $\mathcal{E}=\mathcal{V}\oplus\mathcal{V}_{{}_{\perp}}$, or any vector $w\in\mathcal{E}$ could be written in a unique way as $w=w_{\mathcal{V}}+w_{\mathcal{V}_{\perp}}$, with $w_{\mathcal{V}}\in\mathcal{V},w_{\mathcal{V}_{\perp}}\in\mathcal{V}_{\perp}$. In that case, the map $\mathcal{P}_{{}_{\mathcal{V}}}:w\mapsto w_{\mathcal{V}}$ is the orthogonal projection to $\mathcal{V}$. It satisfies
(2) $\langle\mathcal{P}_{{}_{\mathcal{V}}}w,v\rangle=\langle w,v\rangle\text{ for }v\in\mathcal{V},w\in\mathcal{E}.$